OpenStudy (waheguru):

When WIll This Rocket Hit the Ground y = -5t^2+23t+10 Is 5.4 seconds the correct answer?

OpenStudy (waheguru):

The factored form is h=−(5t+2)(t−5)

OpenStudy (waheguru):

So we subtract the two x-intercepts right, giving 5.4 but the answers in the textbook say 5 seconds so I am confused

geerky42 (geerky42):

Rocket hit the ground when y = 0 so 0 = (5t+2)(t−5)

OpenStudy (waheguru):

Yea, well I get 5.4 by doing that

OpenStudy (anonymous):

it doesn't factor that way

geerky42 (geerky42):

How did you get 5.4? because I get 5.

OpenStudy (anonymous):

i get neither

OpenStudy (waheguru):

How did you guys factor it?

OpenStudy (anonymous):

\[-5t^2+25t+10=0\] divide by \(-5\) get \[t^2-5t-2=0\] the use the quadratic formula

geerky42 (geerky42):

satellite's right, (5t+2)(t−5) is not correct

OpenStudy (unklerhaukus):

-5(5.372281323269014)^2+25(5.372281323269014)+10=0

OpenStudy (waheguru):

We have yet to learn the quadeatic formula, we do this by factoring and looking at the x intercepts

geerky42 (geerky42):

Factor -5 out. -5t^2+25t+10 = -5(t^2 - 5t - 2)

OpenStudy (anonymous):

then either there is a typo, or you have to guess at the solution, because this one does not factor over the integers

OpenStudy (waheguru):

There is a tyop it was actullt 23t

OpenStudy (anonymous):

imagine...

OpenStudy (anonymous):

\[-5t^2+23t+10=0\] or \[5t^2-23t-10=0\]now factor \[(t-5)(5t+2)=0\] so \(t=5\) or \(t=-\frac{2}{5}\)

OpenStudy (anonymous):

pick \(t=5\) because you are not going backwards in time

OpenStudy (waheguru):

@satellite73 say, both the x-intercepts were positive, the what would we doo

OpenStudy (waheguru):

For an example, what if the two x-intercepts were positive like 5 and 1

OpenStudy (unklerhaukus):

if you have two positive intercepts , then when t= the first intercept, the rock will be coming out of the ground and landing when t = the latter intercept ,